OpenStudy (anonymous):

Solutions to csc 3x= 1/9

OpenStudy (anonymous):

\[\csc 3 x=\frac{ 1 }{ 9 },\frac{ 1 }{ \sin 3 x }=\frac{ 1 }{ 9 },\sin 3x=9\] which is impossible as \[\left| \sin 3x \right|\le 1\]